OpenStudy (anonymous):

Please help Find all solutions to the equation. cos2x + 2 cos x + 1 = 0

OpenStudy (helder_edwin):

u have \[ \large \cos^2x+2\cos x+1=0 \] set \(v=\cos x\) then u have \[ \large v^2+2v+1=0 \]

OpenStudy (anonymous):

cos 2x = cos (x+x) = cosxcosx-sinxsinx = cos^2 x - sin ^2 x so your expression is cos^2 (x) - sin^2 (x) + 2 cos (x) + 1 = 0 sin^2 x is just 1- cos^2 (x) to substitute: cos^2 (x) - (1-cos^2 (x)) + 2 cos (x) + 1 = 0 = cos^2 (x) - 1+ cos^2 (x) + 2 cos (x) + 1 = 0 = 2 cos^2 (x) + 2 cos (x) = 0 = 2 cos (x) (cos (x) + 1) = 0 so when cos x is 0 or cos x is -1, the whole expression will equal 0. remember the unit circle. cos is 0 every pi/2 and every 3pi/2. cos is -1 every pi. so your values of x given the domain [0, 2pi] should be pi/2, 3pi/2 and pi.

OpenStudy (anonymous):

Thanks so much man